OpenStudy (anonymous):

write the equation that is the translation of y=|x| left 1 unit and up 2 units. PLEASE SHOW WORK!!

OpenStudy (anonymous):

the function is absolute value of x: IxI for any function of x say f(x) to have it move to the left 1 unit you must have f(x+1) that means wherever you have x you now have (x+1). so you have y=IxI to move it one unit to the left you have y=Ix+1I now to move it 2 units up you must add 2 to your function so you have y=Ix+1I + 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome